The sum of a fraction and its reciprocal is 34/15. Find the fraction.
Answer by stanbon(75887)

The sum of a fraction and its reciprocal is 34/15
---------
EQUATION:
x+(1/x) = 34/15
Add the left side:
(x^2+1)/x = 34/15
Cross-multiply to get:
15x^2+15=34x
15x^2-34x+15=0
------------
x=[34+-sqrt(34^2-4*15*15)]/30
x=[34+-16]/30
x=2 or 18/30= 3/5
----------------------------
Cheers,
Stan H.